Question

Cash Budgeting Retail outlets purchase snowboards from Slopes, Inc., throughout the year. However, in anticipation of late summer and early fall purchases, outlets ramp up inventories from May through August. Outlets are billed when boards are ordered. Invoices are payable within 60 days. From past experience, Slopes' accountant projects 20% of invoices will be paid in the month invoices, 50% will be paid in the following month, and 30% of the invoices will be paid two months after the month of invoice. The average selling price per snowboard is $450. To meet demand, Slopes increases production from April through July, because the snowboards are produced a month prior to their projected sale. Direct materials are purchased in the month of production and are paid for during the following month (terms are payment in full within 30 days of the invoice date). During this period there is no production for inventory, and no materials are purchased for inventory. Direct manufacturing labor and manufacturing overhead are paid monthly. Variable manufacturing overhead is incurred at the rate of $7 per direct manufacturing labor-hour. Variable marketing costs are driven by the number of sales visits. However, there are no sales visits during the months studied. Slopes Inc., also incurs fixed manufacturing overhead costs of $5500 per month and fixed non manufacturing overhead costs of $2500 per month.

Projected Sales

May 80 units

June 120 units

July 200 units

August 100 units

September 60 units

October 40 units

Direct materials and direct manufacturing labor utilization and cost

Wood = 5 units per board, $30 per unit, unit= board feet

Fiberglass = 6 units per board, $5 per unit, unit = yard

Direct manufacturing labor = 5 units per board, $25 per unit, unit = hour

The beginning cash balance for July 1, 2012 is $10000. On October 1, 2011, Slopes had a cash crunch and borrowed $30000 on a 6% one year note with interest payable monthly. The note is due October 1, 2012. Using the information provided, you will need to determine whether Slopes will be in a position to pay off this short term debt on October 1, 2012.

1. Prepare a cash budget for the months of July through September 2012. Show supporting schedules for the calculation of receivables and payables.
Answer #1

A G D E F July August September $10,000 $57,850 $74,300 $90,000 $45,000 $27,000 $100,000 $102,850 $101,300 $42,150 $28,550 $2July 10000 =D15 =SUM(D2:D3) =D36 =D4-D5 August =D6 =E15 =SUM(E2:E3) =E36 =E4-E5 September =E6 =F15 =SUM(F2:F3) =F36 =F4-F5 1